# Effects of rotation and anisotropy on the properties of type-II holographic superconductors

### v1

@@ Abstract
The present work concerns the detailed construction of a holographic model for a type-II s-wave superconductor defined on a 5-dimensional anisotropic rotating black hole. We examine the role of rotation and anisotropy on the properties of the superconductor model focusing on the condensate and the AC conductivity, for which we obtain closed formulas, using both analytical and numerical methods. The results reveal that the rotation is responsible for the appearance of a peak and for introducing an exponentially vanishing behavior in the high-frequency limit of the real component of the AC conductivity. Such a behavior aligns with that observed in high-temperature superconductor models and experiments, where the peak and vanishing behavior result from quasiparticle damping, suggesting a plausible relation between the rotation of a black hole and quasiparticle damping effects in a superconducting material. In addition, we provide a detailed construction of the vortex lattice presented in Herrera-Mendoza:2022whz and study its behavior as a function of an external uniform magnetic field. Once again, it is shown that the vortex lattice can be continuously deformed along with a change in the vortex population by virtue of the magnetic field, providing a promising avenue for holographically modeling the vortex lattice deformations observed in experimental superconducting materials. As an observed experimental effect, we describe both the vortex lattice deformation and the vortex population increment under the action of an external magnetic field in the LiFeAs type-II superconductor. These effects supplement those previously found for the FeSe type-II superconductor studied in Herrera-Mendoza:2022whz .
@@ 1 Introduction
Over the past years, the gauge/gravity duality has emerged as a powerful tool for investigating various phenomena within diverse physical contexts. Originally known as the AdS/CFT correspondence Maldacena:1997re , it establishes a bridge between the behavior of a quantum field theory defined on a flat spacetime and the dynamics of a gravitational theory in a higher-dimensional spacetime. This approach has been particularly useful in studying the properties of strongly coupled quantum systems, such as those found in condensed matter physics and nuclear physics, where traditional analytical techniques fail to provide insight. In this regard, interesting explorations were given in condensed matter physics with a holographic description of field theory superconductors Hartnoll:2008vx and superfluids Herzog:2008he .
These seminal works laid the groundwork for future investigations within holographic superconductivity and superfluidity frameworks. A crucial fact is that holographic superconductors and superfluids can arise in a wide range of gravitational contexts supported by diverse matter fields. This includes gravitational backgrounds whose asymptotic is not only AdS but Lifshitz Taylor:2008tg ; Ayon-Beato:2009rgu ; Ayon-Beato:2010vyw ; Ayon-Beato:2019kmz ; Bravo-Gaete:2020ftn ; Bravo-Gaete:2021kgt , which have proven useful to explore field theory systems with quantum critical regions governed by non-relativistic anisotropic scalings between space and time Kachru:2008yh .
Concretely, holographic models for superconductors and superfluids have been extensively explored using asymptotically AdS backgrounds Hartnoll:2008vx ; Hartnoll:2008kx ; Hartnoll:2009sz ; Herzog:2008he ; Gangopadhyay:2012am ; Gangopadhyay:2012np , including effects of adding a uniform external magnetic field that led to the development of a holographic description of an Abrikosov vortex latticeNakano:2008xc ; Albash:2008eh ; Albash:2009iq ; Albash:2009ix ; Maeda:2009vf ; Montull:2009fe ; Adams:2012pj ; Xia:2019eje ; Srivastav:2023qof . On the other hand, holographic superconductors defined on non-relativistic backgrounds have been approached using asymptotically Lifshitz spacetimes Bu:2012zzb ; Lu:2013tza ; Natsuume:2018yrg ; Zhao:2013pva . Interestingly, and relevant to our present work, the effects of rotation have also been explored in scenarios with AdS Sonner:2009fk ; Lin:2014tza ; Srivastav:2019ixc and Lifshitz asymptotic Herrera-Mendoza:2022whz .
@@
This work aims to develop a detailed and more comprehensive holographic model for a type-II superconductor that is defined on a rotating and anisotropic black hole configuration. By incorporating a rotating background, our primary goal is to investigate the impact of rotation on the properties of a holographic superconductor, with special attention to the condensate phenomena and AC conductivity for which closed expressions were obtained. The results suggest a relationship between the rotation of the black hole and quasiparticle damping effects induced by impurities or defects in a superconducting material. Additionally, we explore the formation of the vortex lattice and its deformation under an external magnetic field, a common feature of type-II superconductors. It is observed how infinitesimal variations of an external magnetic field influence the dynamics of vortices, changing their population and causing continuous deformations in the vortex lattice of type-II holographic superconductors.
@@
The manuscript is organized as follows: in Sec. II we introduce the holographic setup working in the so-called probe limit. Sec. III is devoted to studying the condensation of the scalar field using both analytical and numerical methods, while Sec. IV presents an analytical treatment of the AC conductivity supplemented with numerical computations. In addition, in Sec. V we investigate the effect of a uniform external magnetic field on the vortex lattice, uncovering how the field influences the arrangement of vortices. We end in Sec. VI with a discussion and conclusions relevant to our findings.
@@ 2 The setup
The background Lagrangian $\mathcal{L}_{\text{bg}}$ is defined by the Einstein-Hilbert action with cosmological constant $\lambda$, supplemented by a scalar dilaton field $\varphi$ and a gauge field strength $\mathcal{F}_{\mu\nu}$. Moreover, the matter Lagrangian $\mathcal{L}_{\text{matter}}$ is defined by a gauge field strength $F_{\mu\nu}$ and a complex scalar field $\Psi$, charged by the covariant derivative $D_{\mu}\equiv\nabla_{\mu}-iA_{\mu}$. In addition, we take the limit $q\rightarrow\infty$ wherein the Lagrangian $\mathcal{L}_{\text{bg}}$ can be considered decoupled from $\mathcal{L}_{\text{matter}}$, such that we can neglect the backreaction of $\mathcal{L}_{\text{matter}}$ and work the equations of motion associated to $\mathcal{L}_{\text{bg}}$ in order to look for a concrete gravitational solution with a manifest Lifshitz symmetry. Recall the rotating black hole solution for the $\mathcal{L}_{\text{bg}}$ sector obtained in Herrera-Aguilar:2021top
@@
$\begin{split}&f(r)=1-\quantity(\dfrac{r_{h}}{r})^{z+d},\qquad\qquad\qquad\qquad\quad e^{-b\varphi}=\kappa\dfrac{Q^{2}\ell^{2}}{(z-1)(z+d)}\left(\dfrac{\ell}{r}\right)^{2d},\\ &\mathcal{F}_{tr}=\Xi\dfrac{(z-1)(z+d)}{\kappa Q\ell^{2}}\left(\dfrac{r}{\ell}\right)^{d+z-1},\qquad\mathcal{F}_{\phi_{i}r}=-a_{i}\dfrac{(z-1)(z+d)}{\kappa Q\ell^{2}}\left(\dfrac{r}{\ell}\right)^{d+z-1},\\ &\Lambda=-\dfrac{(z+d)(z+d-1)}{2\ell^{2}},\qquad\qquad\qquad b=2\sqrt{\dfrac{\kappa d}{z-1}},\end{split}$ (4)
where $n=[(d+1)/2]$ is the maximal number of independent rotation planes possible in $d+2$ dimensions, each plane having associated a rotation parameter $a_{i}$, while $\Xi=\sqrt{1+\sum_{i=1}^{n}a_{i}^{2}/\ell^{2}}$, and $d\vec{y}{}^{2}_{(d-n)}=dy_{k}dy^{k}$ is the Euclidean metric in $d-n$ dimensions. We take the particular case of 5 dimensions and assume only one non-trivial rotation parameter ($a_{1}=a$ and $a_{2}=0$). In this case, the metric is reduced to
@@
with $d\vec{x}{}^{2}=dx^{2}+dy^{2}$ and $f(r)=1-\left(\dfrac{r_{h}}{r}\right)^{z+3}$.
@@ 3 Condensation of the scalar field
We know that superconductors are characterized by the presence of a scalar field acting as the order parameter that condensates at some critical temperature. Here, we want to explore whether the scalar field $\Psi$ possesses such a property. To do so, it is sufficient to consider real scalar and gauge fields with the simplest ansätze consistent with the equations (2.1) and (9), i.e.
@@
$\quad\,+\dfrac{\ell^{2}m^{2}}{u^{2}f}\Biggr{\}}\psi=0$ (11c)
Now, we propose Taylor series expansions of the fields around the horizon $u=1$
$\psi(u)$ $=\psi(1)+\psi^{\prime}(1)\quantity(u-1)+\dfrac{\psi^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots$ (12a)
$A_{t}(u)$ $=A_{t}(1)+A_{t}^{\prime}(1)\quantity(u-1)+\dfrac{A_{t}^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots$ (12b)
$A_{\phi}(u)$ $=A_{\phi}(1)+A_{\phi}^{\prime}(1)\quantity(u-1)+\dfrac{A_{\phi}^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots$ (12c)
Here, to have a regular gauge field at the horizon, we need the boundary conditions
$A_{t}(1)=0,\qquad A_{\phi}(1)=0,$ (13)
while the boundary condition for $\psi(1)$ is arbitrary but finite. The remaining coefficients are determined by making the expansions (12) consistent with the field equations (11). After some algebraic computations, we find
Note that the higher-order terms are also expressed in terms of the lowest-order ones.
On the other hand, in order for the fields to have a correct interpretation from the holographic point of view, the near boundary ($u\rightarrow 0$) behavior must have the form $\psi\sim u^{\Delta}$, $A_{t}\sim u^{\alpha_{1}}$ and $A_{\phi}\sim u^{\alpha_{2}}$ with $\Delta$, $\alpha_{1}$ and $\alpha_{2}$ non-negative numbers. After substitution into the field equation (11) and taking the limit $u\rightarrow 0$ we are left with the regular solutions
@@
provided $z\leq 3$, where $\Delta$ is interpreted as the scaling dimension of the dual scalar operator (order parameter) $\mathcal{O}$, which is found to be $\Delta_{\pm}=[(z+3)\pm\sqrt{(z+3)^{2}+4\ell^{2}m^{2}}]/2$. Moreover, $\mu$ and $\nu$ are interpreted as potentials in the dual field theory, while $\rho$ and $\zeta$ are respectively the charge and the current density. For this specific field configuration, these densities are related to each other by $\zeta=-\dfrac{a}{\Xi}\rho$ [^1] Note that there is another branch of solutions, $\alpha_{1}=\alpha_{2}=z+1$ with $\zeta=-\dfrac{\Xi\ell^{2}}{a}\rho$. Nevertheless, they are inconsistent with the upcoming procedure and will be discarded. In addition, to keep the arguments as simple as possible we also impose the boundary condition $J_{-}=0$ and set $J_{+}=J$, $\Delta_{+}=\Delta$, i.e., we are considering $J_{+}$ as the expectation value of the order parameter with scaling dimension $\Delta$, and $J_{-}$ as its source.
Thus far we have found expressions for both regimes, near horizon $(u=1)$ and asymptotic $(u=0)$ solutions. In what follows, we impose the fundamental requirement that there is an intermediate point $u=u_{m}$ in the bulk where the asymptotic solutions, (15) and (12), match and their first derivatives as well. This assumption leads us to four independent relations
From (16a) and (16b) we arrive at
where
Similarly, from (16d) we obtain
Figure 1: The condensation operator for different values of the rotation parameter. Figure (a) portrays the condensation of the dual operator derived through the matching method as shown in equation (3) when $u_{m}=0.8$, $z=2$, and $\ell^{2}m^{2}=-3$. On the other hand, Figure (b) presents the results obtained by numerically solving the system (11) using the shooting method with $z=2$ and $\ell^{2}m^{2}=-3$.
At this point let us recall the temperature of the black hole (5), which has the form Herrera-Aguilar:2021top
$T=\dfrac{1}{4\pi}\dfrac{(z+3)r_{h}^{z}}{\ell^{z+1}\Xi},$ (20)
allowing us to write (19) as an explicit function of the black hole temperature
$\psi(1)^{2}=\dfrac{(z+3)}{2(1-u_{m})}\quantity[\dfrac{\ell^{z}(3-z)}{\gamma\Xi^{1+3/z}}\quantity(\dfrac{z+3}{4\pi\ell^{z+1}})^{3/z}\dfrac{\rho u_{m}^{2-z}}{T^{3/z}}-(z-2)(1-u_{m})-1].$ (21)
The holographic dictionary facilitates a connection between the expectation value of the dual operator $\expectationvalue{\mathcal{O}}$ and the coefficient associated with the subleading order of the scalar field $J$. Concretely we get, $\expectationvalue{\mathcal{O}}=\sqrt{2}r_{h}^{\Delta}J$, where the critical temperature is reached when $\expectationvalue{\mathcal{O}}=0$, or equivalently $\psi(1)=0$. Thus we arrive at the critical temperature for the condensation to occur
Therefore, the expectation value for the condensation operator reads
@@
From this, we identify different behaviors depending on the sign of the power $32\Delta-$ 3. Nevertheless, near the critical point $T=T_{c}$, the behavior is the same for any of the cases.
Figure 1 depicts the condensation of the dual operator $\expectationvalue{\mathcal{O}}$ from two different perspectives: analytical (matching method) and numerical (shooting method). While both methods measure the same quantity, there are significant variations between them, as observed in the plots. However, it is not surprising that they do not yield identical results. Recall that the matching method involves an additional parameter, $u_{m}$, whose selection strongly influences the results for $\expectationvalue{\mathcal{O}}$. Hence, it only offers a qualitative explanation of the phenomenon. In contrast, the shooting method provides a more accurate depiction as it does not introduce any additional parameters and aligns with the anticipated outcome of the matching method. In other words, both figures exhibit the impact of varying the rotation parameter, which only decreases the amplitude of the expectation value for the condensation operator when increasing the rotation.

### v2

# Effects of background rotation and anisotropy in the holographic description of type-II superconductors

@@ Abstract
The present work concerns the detailed construction of a holographic model for a type-II s-wave superconductor defined on a 5-dimensional anisotropic rotating black hole. We examine the role of rotation and anisotropy on the properties of the superconductor model focusing on the condensate and the AC conductivity, for which we obtain closed formulas, using both analytical and numerical methods. The results reveal that the rotation is responsible for the appearance of a peak and for introducing an exponentially vanishing behavior in the high-frequency limit of the real component of the AC conductivity. Such a behavior aligns with that observed in high-temperature superconductor models and experiments, where the peak and vanishing behavior result from quasiparticle damping, suggesting a relation between the rotation of a black hole and quasiparticle damping effects due to impurities or defects in a superconducting material. This relation supplements the holographic dictionary of the gravity/Condensed Matter Theory correspondence. In addition, we provide a detailed construction of the vortex lattice presented in Herrera-Mendoza:2022whz and study its behavior as a function of an external uniform magnetic field. Once again, it is shown that the vortex lattice can be continuously deformed along with a change in the vortex population by virtue of the magnetic field, providing a promising avenue for holographically modeling the vortex lattice deformations observed in experimental studies with superconducting materials. As a concrete example, we describe both the vortex lattice deformation and the increment of the vortex population under the action of an external magnetic field in the LiFeAs type-II superconductor. These effects supplement those previously found for the FeSe type-II superconductor studied in Herrera-Mendoza:2022whz .
@@ 1 Introduction
Over the past years, the gauge/gravity duality has emerged as a powerful tool for investigating various phenomena within diverse physical contexts. Initially introduced as the AdS/CFT correspondence Maldacena:1997re , this duality establishes a relation between a quantum field theory at strong coupling defined on a flat spacetime and a gravitational theory in a higher-dimensional spacetime. This approach has proven particularly useful for studying strongly coupled quantum systems, such as those found in condensed matter physics and nuclear physics, where traditional analytical techniques fail to provide insight. In this regard, notable applications include the holographic description of high temperature superconductors Hartnoll:2008vx and superfluids Herzog:2008he . From the condensed matter theory picture, strong-coupling superconductivity is understood as high-temperature superconductivity induced by a very strong phonon-mediated pairing interaction Rainer_1995 . Within this framework, typical energies of superconducting electrons are similar to energies of lattice excitations that mediate the attractive interaction. Thus, superconductivity is sensitive to the dynamical properties of these excitations, which departs from the laws of weakly-coupled or low-temperature superconductivity. These physical properties motivate the search for dual gravitational descriptions capable of capturing such nontrivial dynamics.
These seminal works laid the groundwork for subsequent developments within the frameworks of holographic superconductivity and superfluidity. Notably, holographic models of superconductors and superfluids can emerge from a variety of gravitational setups supported by diverse matter content. These include not only asymptotically AdS backgrounds but also Lifshitz geometries Taylor:2008tg ; Ayon-Beato:2009rgu ; Ayon-Beato:2010vyw ; Ayon-Beato:2019kmz ; Bravo-Gaete:2020ftn ; Bravo-Gaete:2021kgt , which have proven valuable for modeling field theory systems with quantum critical behavior governed by non-relativistic anisotropic scaling between space and time Kachru:2008yh . More concretely, holographic models of superconductors and superfluids have been extensively developed in asymptotically AdS backgrounds Hartnoll:2008vx ; Hartnoll:2008kx ; Hartnoll:2009sz ; Herzog:2008he ; Gangopadhyay:2012am ; Gangopadhyay:2012np . These studies include the effects of uniform external magnetic fields, which led to the holographic modeling of Abrikosov vortex lattices Nakano:2008xc ; Albash:2008eh ; Albash:2009iq ; Albash:2009ix ; Maeda:2009vf ; Montull:2009fe ; Adams:2012pj ; Xia:2019eje ; Srivastav:2023qof . On the other hand, non-relativistic holographic superconductors have been investigated in the context of asymptotically Lifshitz spacetimes Bu:2012zzb ; Lu:2013tza ; Natsuume:2018yrg ; Zhao:2013pva . Interestingly, and of particular relevance to the present work, rotational effects have also been studied in both AdS Sonner:2009fk ; Lin:2014tza ; Srivastav:2019ixc and Lifshitz scenarios Herrera-Mendoza:2022whz .
@@
This work develops a comprehensive holographic model of a type-II superconductor defined on a rotating, anisotropic black hole background. By incorporating a rotating background, our primary goal is to investigate the impact of rotation on the properties of a holographic superconductor, with special attention to the condensate phenomena and AC conductivity for which closed expressions were obtained. The results indicate the following relationship
[…]
Additionally, we explore the formation of the vortex lattice and its deformation under an external magnetic field, a common feature of type-II superconductors. It is observed how infinitesimal variations of an external magnetic field influence the dynamics of vortices, changing their population and causing continuous deformations in the vortex lattice of type-II holographic superconductors. Remarkably, both of these physical effects are experimentally observed in the LiFeAs type-II superconductor under the action of an external magnetic field.
@@
The manuscript is organized as follows: in Sec. II we introduce the holographic setup working in the so-called probe limit. Sec. III is devoted to studying the condensation of the scalar field using both analytical and numerical methods, while Sec. IV presents an analytical treatment of the AC conductivity supplemented with numerical computations. In addition, in Sec. V we investigate the effect of a uniform external magnetic field on the vortex lattice, uncovering how the field influences the arrangement of vortices. We conclude in Sec. VI with a discussion of our relevant findings and their implications.
@@ 2 The setup
The background Lagrangian $\mathcal{L}_{\text{bg}}$ is defined by the Einstein-Hilbert action with cosmological constant $\lambda$, supplemented by a scalar dilaton field $\varphi$ and a gauge field strength $\mathcal{F}_{\mu\nu}$. While the matter Lagrangian $\mathcal{L}_{\text{matter}}$ is defined by a gauge field strength $F_{\mu\nu}$ and a complex scalar field $\Psi$, charged by the covariant derivative $D_{\mu}\equiv\nabla_{\mu}-iA_{\mu}$. In addition, we take the limit $q\rightarrow\infty$ wherein the Lagrangian $\mathcal{L}_{\text{bg}}$ can be considered decoupled from $\mathcal{L}_{\text{matter}}$, such that we can neglect the backreaction of $\mathcal{L}_{\text{matter}}$ and work the equations of motion associated to $\mathcal{L}_{\text{bg}}$ in order to look for a concrete gravitational solution with a manifest Lifshitz symmetry.
Recall the rotating, anisotropic black hole solution for the $\mathcal{L}_{\text{bg}}$ sector obtained in Herrera-Aguilar:2021top
@@
$\begin{split}&f(r)=1-\quantity(\dfrac{r_{h}}{r})^{z+d},\qquad\qquad\qquad\qquad\quad e^{-b\varphi}=\kappa\dfrac{Q^{2}\ell^{2}}{(z-1)(z+d)}\left(\dfrac{\ell}{r}\right)^{2d},\\ &\mathcal{F}_{tr}=\Xi\dfrac{(z-1)(z+d)}{\kappa Q\ell^{2}}\left(\dfrac{r}{\ell}\right)^{d+z-1},\qquad\mathcal{F}_{\phi_{i}r}=-a_{i}\dfrac{(z-1)(z+d)}{\kappa Q\ell^{2}}\left(\dfrac{r}{\ell}\right)^{d+z-1},\\ &\Lambda=-\dfrac{(z+d)(z+d-1)}{2\ell^{2}},\qquad\qquad\qquad b=2\sqrt{\dfrac{\kappa d}{z-1}}.\end{split}$ (4)
This gravitational configuration describes a rotating Lifshitz-like black hole in $d+2$ dimensions, supported by a dilaton field, a Maxwell field with electric and magnetic components, and a negative cosmological constant. It includes up to $n=[(d+1)/2]$ independent rotation planes, each associated with a parameter $a_{i}$, leading to anisotropy and rotational effects encoded in the metric through $\Xi=\sqrt{1+\sum_{i=1}^{n}a_{i}^{2}/\ell^{2}}$. The spacetime exhibits Lifshitz scaling with dynamical exponent $z$, a black hole horizon at $r=r_{h}$, and flat transverse directions described by the Euclidean metric $d\vec{y}{}^{2}_{(d-n)}=dy_{k}dy^{k}$, making it a suitable background for holographic models of strongly correlated systems with anisotropic and rotating features. The temperature $T$ and angular momentum $\mathcal{J}_{i}$ of this configuration are respectively given by
$\quad T=\dfrac{1}{4\pi}\dfrac{(z+d-2)r_{h}^{z}}{\ell^{z+1}\Xi},\quad\quad\mathcal{J}_{i}=\dfrac{a_{i}\Xi(d-z)V_{(d-2)}r_{h}^{z+d-2}}{2\kappa\ell^{z+[d/2]}}.$ (5)
See Herrera-Aguilar:2021top for a deeper analysis concerning the physical properties of this background.
For the purposes of this work, we consider a particular case of spacetime (2) in five dimensions, assuming only one non-trivial rotation parameter ($a_{1}=a$, $a_{2}=0$). In this case, the metric reduces to
@@
with $d\vec{y}{}^{2}=dx^{2}+dy^{2}$ and $f(r)=1-\left(\dfrac{r_{h}}{r}\right)^{z+3}$.
@@ 3 Condensation of the scalar field
Holographic superconductivity is characterized by the presence of a scalar field acting as the order parameter that condensates below some critical temperature. Here, we want to explore the condensation of the scalar field $\Psi$ in our field configuration. To achieve this, we consider real scalar and gauge fields with the simplest ansätze consistent with the equations (2.1) and (10), i.e.
@@
$\quad\,+\dfrac{\ell^{2}m^{2}}{u^{2}f}\Biggr{\}}\psi=0.$ (12c)
In order to explore the condensation phenomenon, we need to solve this system at the asymptotic boundary ($u=0$). In this regard, consistent with the holographic prescription, it is simple to show that the system adopts the following boundary solutions
@@
provided $z\leq 3$, where $J_{-}$ ($J_{+}$) can be identified as the source (the expectation value) of the scalar operator $\mathcal{O}$, with $\Delta_{\pm}=[(z+3)\pm\sqrt{(z+3)^{2}+4\ell^{2}m^{2}}]/2$ the corresponding scaling dimensions. To ensure the existence of normalizable modes near the asymptotic boundary, the mass of the scalar field must satisfy the Breitenlohner-Freedman (BF) bound for massive scalars in a Lifshitz background, given by $m^{2}\ell^{2}\geq-\frac{(z+3)^{2}}{4}$. Additionally, $\mu$ and $\nu$ are interpreted as potentials in the dual field theory, while $\rho$ and $\zeta$ represent the charge density and the current density, respectively. In this specific field configuration, these densities are related by $\zeta=-\dfrac{a}{\Xi}\rho$.
The holographic dictionary facilitates the relation $\expectationvalue{\mathcal{O}}=\sqrt{2}r_{h}^{\Delta}J$, in which $J$ is a function of the background parameters.
Figure 1: Profiles of the condensation operator taking different values of the rotation parameter and considering the boundary condition $J_{+}=0$, $J_{-}\sim\expectationvalue{\mathcal{O}_{1}}$, and fixed parameters $\ell=1$, $z=2$ and $\ell^{2}m^{2}=-4$. The left panel shows the condensation of the dual operator derived via the matching method (equation (3)), taking $u_{m}=0.8$. The right panel displays numerical results from solving (12) using the shooting method.
In Appendix A, we present a detailed analytic derivation for $J$ (and $\expectationvalue{\mathcal{O}}$) in terms of the black hole parameters, using the matching method. This leads to the following expression for the expectation value of the order parameter:
@@
Here, $u_{m}$ is an intermediate bulk point introduced by the matching method, where the near-horizon and boundary expansions are smoothly joined. Moreover, the critical temperature $T_{c}$ is determined by imposing the condition $\expectationvalue{\mathcal{O}}=0$, yielding
The expression (3) can be viewed as a holographic generalization of the Ginzburg–Landau order parameter when $z\to 3$ and $\Delta\to 3/2$, and it reproduces the characteristic scaling of the BCS gap near $T_{c}$ for $z\to 1$ with the same scaling dimension. For general values of the dynamical exponent $z$ and the scaling dimension $\Delta$, it describes novel critical regimes that go beyond conventional superconducting models.
Figure 1 depicts the condensation of the dual operator $\expectationvalue{\mathcal{O}}$ from two different perspectives: analytical (matching method) and numerical (shooting method). While both methods measure the same quantity, there are significant variations between them, as observed in the plots. However, it is not surprising that they do not yield identical results. Recall that the matching method involves an additional parameter, $u_{m}$, whose selection strongly influences the results for $\expectationvalue{\mathcal{O}}$. Hence, it only offers a qualitative explanation of the phenomenon. In contrast, the shooting method provides a more accurate depiction as it does not introduce any additional parameters and aligns with the anticipated outcome of the matching method. In other words, both figures exhibit the impact of varying the rotation parameter, which only scales the amplitude of the expectation value for the condensation operator as the rotation increases.